1. The following table summarize 200 endothermic reactions involving sodium bicarbonate. Heat Absorbed (cal) Below Target (B) Above target (B) Final Temperature Conditions Final Temperature 266 K Final Temperature .271 K Final Temperature 274 K 12 42 54 40 16 36 Note that the total number of reactions in this table is 12 + 40 +42 + 16 +54 + 36 -200, and below we consider two different eventS Let A denote the event that a reactions final temperature is 271K or less (i.e., the first two rows). Let B denote the event that the heat absorbed in a reaction is below target (i.e., the first column) (i) Determine the number of reactions in each of the following events. (a) A (b) B (c) An B (d) A (e) AUB (f) AUB(g) An B (ii) Determine the following probabilities. (a) P(A) (b) P(B) (c) P(AnB) (d) P(A) (e) P(AU B) (f) P(AUB) (g) P(AnB)
Answer #1

i)

(a)

Number of reactions in A is

12+40+42+16=110

(b)

Number of reactions in B is

12+42+54=108

(c)

Number of reactions in A \cap B is

12+42=54

(d)

Number of reactions in A` is

200-110=90

(e)

Number of reactions in A \cup B is

A \cup B= A + B - A \cap B

=110+108-54= 164

(f)

Number of reactions in A \cup B` is

A \cup B`= A + B` - A \cap B`

=110+92-56= 146

(g)

Number of reactions in A` \cap B` is

=36

(ii)

(a)

The probability of A is

110/200=0.55

(b)

The probability of B is

108/200=0.54

(c)

The probability of A \cap B is

54/200=0.27

(d)

The probability of A` is

90/200=0.45

(e)

The probability of A \cup B is

164/200=0.82

(f)

The probability of A \cup B` is

146/200=0.73

(g)

The probability of  A` \cap B` is

=36/200=0.18